Shutdown Activities

From NPDGamma Wiki

Jump to: navigation, search

Jan-2014 Shut Down Activities and when Start LH2 Filling

  1. Check M1 noise and HV
  2. Install low efficient counting monitor
    1. we still keep M4 behind the counting monitor, check signal and HV
    2. check that the scaler is functioning and see that data is going to the DAQ
    3. check counting detector
  3. LH2 Target
    1. get scroll pump fixed
    2. get the TMP151 fixed and back to vacuum line
    3. a plan for OPC
      1. - check thermometers - offsets
      2. - fix the intermittent thermometer #11?
      3. - start pumping OPC
  4. When start filling
    1. beam back on Feb-10th at 8am
    2. take data with empty target
    3. start filling and monitoring of transmission
    4. after fill follow the conversion and change conversion rate
    5. back to taking data as soon as possible
    6. since power outage we need do polarimetry

Updated July-07 Gantt Chart for our Shut Down Activities

http://battlestar.phys.utk.edu/~target/NPDGammaShutDownActivities.png Our activity list during shut down]


  1. Monitors - a plan is to use only with M1 and use M2 for diagnostics
    • HV from 326V batteries
    1. we have batteries boxed with connectors but
    2. we need three short BNC twinax cables
    • M4 closer to target
    1. needs a no vibrating stand
    2. stand should also allow mounting of preamp
    3. signal cable routing
    4. preamp powering from VME3 rack
    5. build better neutron shielding for room background
    6. fix the malfunctioning Artfex preamp - in process
  2. SF - ramp, switcher, monitoring seem to work - no work planned here but
    • we need to look at the additional shielding of the RF power cable
    • also we need to check if we should move the LED powering to other source than VME3 rack
    • we have also to measure rf power leakage out from aluminum SF container, we can add extra Al to areas that beam is not interacting with
    • we also need to check that there is no rf leakage around the SF rack - like rf power from the donate
    • check cable routing - signal - power - SF signal and RF power
  3. Magnetic field
    • get our Danfysik power supply fixed
    • test Indiana's power supply
  4. Detector
    • one detector #xx seems to be little unstable, do we need to open the preamp and maybe set little tighter freq cut off
    • batteries are always our concern - do we need open a few of them and measure voltages
    • add extra neutron shielding outside
    • study to adding more neutron shielding inside - requires removal of SF - space between detector and cryostat is tight
  5. Electronics
    • get three other Bessel filter boards built, mounted and tested
    • VME 3 problem
    • clean cabling in racks
    • remove HV NIM bin out
    • fix the broken fiber cables
  6. On/off line analysis routines
  7. LH2 target
    • keep pumping vacuum
    • leak check for next cool down
    • He channel flow
    • test the failed compressor - by mounting back to refrigerator
    • other routine maintenance ...
    • fix problems with Target Signal Processing System by Melanie
    • water flow monitor and control valve for turbo pump
  8. Do we need to add extra shielding for the experiment to reduce room background
  9. Beam collimators
    • check that we have right collimators available
    • improve the mounting of the collimators on the led wall
  10. Fix shutter controller

Schedule for Summer 2012 Tasks

Schedule is driven by the following main tasks:

  1. Effort to fix VME3, when done we can start the noise asymmetry measurements if the Bessel filter circuits have been installed and tested.
  2. Requirement to have beam on the empty target - this will delay the production by a week.
Personal tools